# A Thorough Assessment of the Non-IID Data Impact in Federated Learning

### v1

@@ V-A2 Convergence
Highlight 5: The higher the non-IIDness level of labels, the more required rounds to achieve convergence.
Table IV: The number of rounds required for FedAvg, Rand, FedProx, POC, and MOON to achieve 90% of the maximum accuracy reached in different levels of non-IID cases as determined by HD over CIFAR10 for K = 30.
@@
Table IV examines the algorithms’ convergence from a different perspective. We investigate the performance of each aggregation approach across a certain level of non-IID scenario independent from the others. For each specific non-IID situation described by HD, we determine how many rounds each aggregation algorithm requires to achieve 90% of its maximum accuracy. Therefore, regardless of the aggregation algorithm, as the non-IIDness of the data partitioning over the clients increases, more communication rounds are necessary to reach a convergence point (where the accuracy gets stable). We observe the latter because the data distributions among clients are sufficiently dissimilar, so the models built for each client are only optimal about their data, which diverges the model from the optimal case. As the training progresses, the weights delivered by the server to the clients improve because they get optimized by considering all of the data across all clients.
Table V: Mean and standard deviation Accuracy for each dataset for CL, FedAvg, Rand, FedProx, Power-Of-Choice, and MOON, considering different levels of feature skewness measured by FHD for K = 30. Each model has undergone ten different trials.
@@
V-B Feature Skew
This section examines how feature skewness in the client data affects the models’ performance.
V-B1 Classification power
In this subsection, we focus on the findings from the simulations to compare different aggregation techniques and datasets regarding their classification power (a.k.a accuracy) in the presence of feature skew over the clients’ data. Table V summarizes the models’ performance derived from different aggregation algorithms under varying degrees of non-IID feature distributions, as indicated by FHD.
Highlight 6: The performance of FL-generated models is lower than CL-generated models.
@@ V-B1 Classification power
Highlight 7: The model’s performance in image datasets remains unaffected by increasing the level of feature non-IIDness.
Consider only the image datasets (CIFAR10, FMNIST) and the models generated in FL. Regardless of the aggregation algorithm employed, the performance of the final model remains stable across different levels of feature non-IIDness, consistently converging to specific values for each aggregation method. The latter pattern could occur due to convolutional layers in image analysis, where alterations in specific pixel values have minimal impact on the outcome of the convolution process in the Gaussian-noise method.
@@
Having tabular datasets (Covtype, Physionet) and using the Hist-Dirichlet approach shows that increasing the degree of feature non-IIDness does not impact the performance. However, when dealing with Gaussian noise, if we surpass FHD=0.9, there’s a noticeable decline in performance. This decline occurs because the data become so dissimilar (too noisy) from each other that even within the CL learning, the model struggles to learn from the data.
@@
Highlight 9: In scenarios where the features are non-IID partitioned across clients, MOON demonstrates superior performance compared to all other aggregation algorithms for tabular datasets.
@@
Table V validates that no particular algorithm outperforms others in image datasets, as they yield comparable final performance metrics. MOON emerges as the top-performing algorithm in both versions in tabular datasets, surpassing all other algorithms. Its performance is nearly equivalent to that of models trained in CL. The latter occurs since MOON can immediately start learning meaningful contrasts between differences in the label using the provided features of the tabular dataset. On the other hand, for images, the model first needs to learn to extract meaningful features from raw pixels before it can start contrasting different object classes effectively. For example, consider the Physionet dataset, which contains features such as age, sex, heart rate, and P-R interval. In that case, each feature has a clear medical interpretation, and the model can directly use the mentioned feature values without learning initial representations. Conversely, for CIFAR10, the model must learn to extract meaningful features from raw pixels before contrastive learning can be effective.
V-C Quantity Skew
@@
This subsection examines how quantity skewness in the client data affects the models’ performance.
@@
Table VI: Mean and standard deviation Accuracy for each dataset for CL, FedAvg, Rand, FedProx, Power-Of-Choice, and MOON, considering different levels of non-IID partitioning of the record quantities measured by QHD for $K=30$. Each model has undergone five different trials.
@@
V-C1 Convergence
In this subsection, we concentrate on the results of the simulations, aiming to contrast various aggregation methods and datasets in terms of their convergence.
Highlight 10: Feature skew doesn’t alter model convergence point
Table VII: The number of rounds required for FedAvg, Rand, FedProx, POC, and MOON to achieve 90% of the maximum accuracy reached in different levels of feature non-IID cases as determined by FHD over CIFAR10 and Covtype using Hist Dirichlet for K = 30
Feature distribution skew Hist Dirichlet CIFAR10 FedAvg 5 5 5 5 4
Rand 5 5 5 5 4
FedProx 5 5 5 5 4
POC 5 5 5 5 4
MOON 8 7 7 8 7
Covtype FedAvg 3 3 3 3 4
Rand 3 3 3 3 4
FedProx 3 3 3 3 4
POC 3 3 3 3 4
MOON 4 4 4 4 5
Table VII presents an alternative viewpoint on the latter concept. It outlines the iterations needed for FedAvg, Rand, FedProx, POC, and MOON to achieve 90% of the maximum accuracy across various degrees of feature non-IID conditions, as characterized by FHD, using the CIFAR10 dataset. Increasing the non-IIDness of features within the data has minimal impact on the model’s ability to converge to its optimal performance.
V-C2 Classification power:
In the following paragraphs, we concentrate on the simulation results to evaluate various aggregation methods and datasets regarding their classification accuracy, particularly considering the impact of quantity skew on the clients’ data.
Highlight 11: The presence of quantity skewness in the client’s data does not affect the performance of the final model.
Examining Table VI and considering each aggregation algorithm separately, it is evident that the performance of the final models remains consistent across various levels of quantity skewness in the clients’ records. This phenomenon occurs regardless of the chosen aggregation algorithm, confirming that quantity skewness does not affect model performance.
V-C3 Convergence
In this subsection, we focus on the findings obtained when different aggregation algorithms regarding their learning process and the smoothness of training.
@@ V-C3 Convergence
Table VIII: The number of rounds required for FedAvg, Rand, FedProx, POC, and MOON to achieve 90% of the maximum accuracy reached in different levels of quantity non-IID cases as determined by QHD over CIFAR10 and Covtype using Min-size Dirichlet method for K = 30
@@
Looking at Table VIII, it is clear that regardless of the degree of non-IIDness in the number of records from each label across clients, all aggregation algorithms converge after a consistent number of communication rounds. This phenomenon occurs due to the similarity in data distributions among clients, where each client’s dataset mirrors others, making any information obtained from all clients also obtainable from just one client.
V-D Spatiotemporal Skew
@@
Table IX: Mean and standard deviation Accuracy for each dataset for CL, FedAvg, Rand, FedProx, Power-Of-Choice, and MOON, considering different levels of non-IID partitioning of the records based on their space (SP) and time (TMP) measured by STHD for $K=30$. Each model has undergone five trials.
@@
Number of times that performed the best 0 0 0 3 7
Table X: Number of rounds required to achieve 90% of the maximum accuracy reached accuracy levels of spatiotemporal non-IID cases as determined by STHD for K = 30
@@
V-D1 Classification power
In this subsection, we focus on the simulation results to evaluate various aggregation methods and datasets regarding classification accuracy. We pay particular attention to the impact of different levels of spatiotemporal skewness among the clients’ data.
Highlight 13: The higher the non-IIDness level in time and space, the worse the model’s performance.
Table IX demonstrates that irrespective of the aggregation algorithm employed, model performance deteriorates when data distribution among clients varies concerning time or space. This phenomenon occurs because increasing the differences among clients’ data based on time and location also raises the level of non-IIDness in the clients’ label distributions. This behavior is evident in Table XI, which displays the HD in label distributions at varying levels of STHD among clients. We also concluded in the Label Skew study section that higher levels of non-IIDness among clients’ data distributions negatively impact the performance of the final model.
V-D2 Convergence
In this subsection, we concentrate on our results related to the models’ convergence when there are variations in the data concerning time and location, considering the results obtained on Serengeti and 5GNTF datasets.
Highlight 14: The higher the non-IIDness level in time and space, the more required rounds to achieve convergence.
As shown in Table X, irrespective of the aggregation algorithm used for training, an increase in the distance between data distributions in terms of time and location results in a need for more rounds to achieve convergence. This trend is evident in both the Serengeti and 5GNTF datasets. When comparing the required rounds for STHD=0 and STHD=90, it is clear that in the worst case (STHD=90), the number of rounds needed is at least twice as high. This observation aligns with our earlier findings on classification performance, specifically from the Label Skew study, where we determined that higher non-IIDness among clients’ label distributions leads to increased rounds required to achieve convergence.
Table XI: HD among clients’ label distributions at varying levels of non-IID partitioning by time and space
Dataset HD = 0 HD = 0.25 HD = 0.50 HD = 0.75 HD = 0.9
Serengeti 0.01 0.09 0.22 0.36 0.53
5GNTF 0.03 0.20 0.29 0.30 0.49
V-E General results
@@
In this section, we provide highlights summarizing the overall results obtained from our experiments, combining the behavior shown before for label, feature, quantity, and spatiotemporal skewness.
Highlight 15: FedProx, POC, and MOON demonstrate superior performance compared to FedAvg and Rand in general FL scenarios
@@
Table XII: The number of cases in which each specific algorithm achieved the best performance for each study
@@
Spatio Temporal Skew 10 0 0 0 3 7
Total best performance 8 3 18 18 31
Table XII summarizes the cases in which each specific algorithm exhibited the best performance compared to other aggregation algorithms across four skewness types considered in our study. In most cases, the FedProx, POC, and MOON aggregation algorithms achieved the best performance, outperforming the simpler FedAvg and Rand algorithms.
Highlight 16: FedProx is more effective on image datasets, while MOON performs better with tabular datasets.
Table XIII: The number of cases in which each aggregation algorithm achieved the best performance for each type of dataset
@@
Image 34 7 3 15 9 0
Tabular 44 1 0 3 9 31
Total best performance 8 3 18 18 31
Table XIII examines the best-performing aggregation algorithms from the perspective of the dataset type used for training. It shows that FedProx outperforms all other algorithms on image datasets in fifteen out of thirty-four cases. In comparison, MOON generally surpasses other algorithms on tabular datasets in thirty out of forty-four cases.
VI Design Insights and Opportunities
We provide some design insights and opportunities intending to help researchers direct their efforts toward solving the effects of data heterogeneity.
Quantifying non-IIDness.
Several works claim that the data non-IIDness affects the performance of FL models [27, 29, 16]. Nevertheless, for the first time, we demonstrate that the effect of the non-IIDness is not the same under all the levels of heterogeneity (see Fig. 4). Therefore, it is vital to quantify the non-IIDness level in FL. This work uses the HD metric to measure the data heterogeneity. However, we encourage researchers to test different metrics such as Jensen-Shannon distance [33], Earth mover’s distance [9], and Total Variation distance [3], among others.
Better methods to tackle high non-IIDness,
This work showcases how the state-of-the-art methods to tackle non-IIDness (Rand, POC, FedProx, MOON) perform against FedAvg. The conclusion is that no algorithm works better than FedAvg in all the scenarios. Moreover, the methods that work better do not show a vast improvement regarding FedAvg under high non-IID scenarios as their gain is at most two percent points. This phenomenon has also been studied and claimed in the scarce work of empirical analysis of non-IID data and methodologies [43, 1, 31, 22]. Therefore, creating methods to appropriately alleviate the effect of high data heterogeneity is needed to evolve and preserve FL. The latter aligns with the open problems reported by Kairouz et al. [18].
Focusing on highly unbalanced data,
In our experiments, we claim that the more significant decrease in performance comparing CL and FL occurs in unbalanced datasets since it relates to the challenge of learning from highly skewed and less representative data (see Table II). Thus, it is relevant to create solutions to tackle non-IIDness by considering the degree of unbalancedness that the labels might have across the clients.

### v2

@@ 5.3 Convergence
Highlight 5: The higher the non-IID data level of labels, the more rounds are required to achieve convergence [38].
Table 7: RTA for FedAvg, Rand, FedProx, POC, and MOON reached in different levels of non-IID cases as determined by HD over CIFAR10 for K = 30.
@@
Table 7 examines the algorithms’ convergence from a different perspective. We investigate the performance of each aggregation approach across a certain level of non-IID scenario, independent of the others. For each specific non-IID situation described by HD, we determine how many rounds each aggregation algorithm requires to achieve 90% of its maximum accuracy. Therefore, regardless of the aggregation algorithm, as the non-IID data partitioning over the clients increases, more communication rounds are necessary to reach a convergence point (where the accuracy gets stable). Such behavior aligns with the findings of Li et al. [38].
We observe the mentioned behavior because the data distributions among clients are sufficiently dissimilar, so the models built for each client are only optimal for their data, which diverges from the optimal case. As the training progresses, the weights delivered by the server to the clients improve because they get optimized by considering all of the data across all clients.
Table 8: Mean and standard deviation Accuracy for each dataset for CL, FedAvg, Rand, FedProx, Power-Of-Choice, and MOON, considering different levels of feature skewness measured by FHD for K = 30. Each model has undergone ten different trials (random seeds).
@@
6 Feature Skew Results
This section examines how feature skew in the client data affects the models’ performance.
6.1 Synthetic Partitioning Method
For simulating feature skew, we employed two diverse methods from FedArtML [33] to test their properties:
Hist-Dirichlet-based method: The process starts by characterizing the attributes of each client using other average values and then subjecting them to a binning procedure. Subsequently, it establishes the participation of each feature category within each client using the DD with a specified $\alpha$. Unlike the Gaussian Noise approach, this method distributes the data among the clients without modifying the features. We measure the non-IID data with the HD among the features across clients (FHD) within the range $\{0,0.25,0.5,0.75,0.9\}$.
6.2 Classification Power
In this subsection, we focus on the findings from the simulations to compare different aggregation techniques and datasets regarding their classification power (a.k.a. accuracy) in the presence of feature skew over the clients’ data. Table 8 summarizes the models’ performance derived from different aggregation algorithms under varying degrees of non-IID feature distributions, as indicated by FHD.
Highlight 6: The performance of FL-generated models is lower than that of CL-generated models.
@@ 6.2 Classification Power
Highlight 7: The model’s performance in image datasets remains unaffected by increasing the feature non-IID data [38].
Consider only the image datasets (CIFAR10, FMNIST) and the models generated in FL. Regardless of the aggregation algorithm employed, the performance of the final model remains stable across different levels of feature non-IID data, consistently converging to specific values for each aggregation algorithm. This behavior is consistent with the observations reported by Li et al. [38].
Such a pattern arises from the robustness of convolutional layers, which extract spatial features while suppressing minor pixel variations. In the Gaussian-noise method, small perturbations do not significantly alter key patterns, as convolutional filters average out noise. Deeper layers further aggregate features, preserving essential information and minimizing the impact on performance.
@@
Having tabular datasets (Covtype, Physionet) and using the Hist-Dirichlet approach shows that increasing the degree of feature non-IID data does not impact the performance. However, when dealing with Gaussian noise, if we surpass FHD=0.9, there’s a noticeable decline in performance.
This decline occurs because the data becomes highly dissimilar and noisy, making it difficult for the model to extract meaningful patterns. Even in CL, where data is typically more stable, excessive noise disrupts feature learning, reducing the model’s generalization ability and leading to performance degradation.
@@
Highlight 9: In scenarios where the features are non-IID partitioned across clients, MOON performs better than all other aggregation algorithms for tabular datasets.
Table 8 validates that no particular algorithm outperforms others in image datasets, as they yield comparable final performance metrics. MOON emerges as the top-performing algorithm in tabular datasets, surpassing all other algorithms. Its performance is nearly equivalent to that of models trained in CL.
@@
This occurs since MOON can immediately start learning meaningful contrasts between differences in the label using the provided features of the tabular dataset. On the other hand, for images, the model first needs to learn to extract meaningful features from raw pixels before it can start contrasting different object classes effectively. For example, consider the Physionet dataset, which contains features such as age, sex, heart rate, and P-R interval. In that case, each feature has a clear medical interpretation, and the model can directly use the mentioned feature values without learning initial representations. Conversely, for CIFAR10, the model must learn to extract meaningful features from raw pixels before contrastive learning can be effective.
6.3 Convergence
In this subsection, we concentrate on the results of the simulations, aiming to contrast various aggregation algorithms and datasets in terms of their convergence.
Highlight 10: Feature skew doesn’t alter the model convergence point.
Table 9: RTA for FedAvg, Rand, FedProx, POC, and MOON reached in different levels of feature non-IID cases as determined by FHD over CIFAR10 and Covtype using Hist Dirichlet for K = 30
Feature distribution skew Hist Dirichlet CIFAR10 FedAvg 5 5 5 5 4
Rand 5 5 5 5 4
FedProx 5 5 5 5 4
POC 5 5 5 5 4
MOON 8 7 7 8 7
Covtype FedAvg 3 3 3 3 4
Rand 3 3 3 3 4
FedProx 3 3 3 3 4
POC 3 3 3 3 4
MOON 4 4 4 4 5
Table 9 presents an alternative viewpoint on the previous highlight. It outlines the iterations needed for FedAvg, Rand, FedProx, POC, and MOON to achieve 90% of the maximum accuracy across various degrees of feature non-IID conditions, as characterized by FHD, using the CIFAR10 dataset. Increasing the non-IID data of features within the data has minimal impact on the model’s ability to converge to its optimal performance.
The minimal impact of feature skew on convergence suggests that while feature distributions differ across clients, the underlying task remains learnable. Unlike label skew, which directly affects class representation in local updates, feature skew primarily alters input variations without disrupting the overall decision boundary. As a result, the global model can still generalize effectively across clients, leading to similar convergence behavior regardless of the degree of feature non-IID data.
7 Quantity Skew Results
@@
This section examines how the quantity skew in the client data affects the models’ performance.
7.1 Synthetic Partitioning Method
We use the MinSize-Dirichlet method included in the FedArtML [33] tool, which specifies the DD’s $\alpha$ value and generates the desired participation proportions for each client. Subsequently, a minimum required size, referred to as ”the minimum number of examples,” is established for each client. Thus, the minimum proportion size, denoted as $MinSize$, is calculated as $\textit{MinSize}=\frac{\textit{MinRequiredSize}}{n}$, where $n$ represents the total number of examples in the centralized dataset. If the designated proportions fall below MinSize, it substitutes them with MinSize. Finally, the proportions are normalized to fall from 0 to 1.
We assess the level of non-IID data using the HD for quantity skew (QHD) within the range $\{0,0.10,0.17\}$. This small range arises because the finite size of the dataset constrains quantity skew. Unlike other skews, the proportions derived from the quantity distribution cannot exhibit extreme divergence, as the total number of samples restricts how unevenly clients can receive data.
@@ 7.1 Synthetic Partitioning Method
Table 10: Mean and standard deviation Accuracy for each dataset for CL, FedAvg, Rand, FedProx, Power-Of-Choice, and MOON, considering different levels of non-IID partitioning of the record quantities measured by QHD for $K=30$. Each model has undergone five different trials.
@@
7.2 Classification Power:
In the following paragraphs, we concentrate on the simulation results to evaluate various aggregation algorithms and datasets regarding their classification accuracy, particularly considering the impact of quantity skew on the clients’ data.
Highlight 11: The quantity skew in the client’s data does not affect the performance of the final model [38].
Table 10 depicts the performance of each aggregation algorithm and dataset, using various levels of non-IID for quantity skew. Examining this table and considering each aggregation algorithm separately, it is evident that the performance of the final models remains consistent across various levels of quantity skewness in the clients’ records. This phenomenon occurs regardless of the chosen aggregation algorithm, confirming that quantity skewness does not affect model performance. This behavior pattern agrees with the results documented by Li et al. [38].
The invariance of model performance to quantity skew suggests that FL aggregation algorithms effectively balance updates regardless of varying client sample sizes. Since clients contribute proportionally to the global model, those with fewer samples still provide functional gradients without disproportionately influencing training. Additionally, standard optimization techniques, such as weighted averaging, mitigate potential biases from data imbalance, ensuring stable performance across different levels of quantity skewness.
7.3 Convergence
In this subsection, we focus on the findings obtained when different aggregation algorithms are considered regarding their learning process and the smoothness of training.
@@ 7.3 Convergence
Table 11: RTA for FedAvg, Rand, FedProx, POC, and MOON reached in different levels of quantity non-IID cases as determined by QHD over CIFAR10 and Covtype using Min-size Dirichlet method for $K=30$
@@
Table 11 shows the RTA for the aggregation algorithms on various levels of quantity non-IID cases for the analyzed datasets. Looking at this table, it is clear that regardless of the degree of non-IID data in the number of records from each label across clients, all aggregation algorithms converge after a consistent number of communication rounds.
The consistent convergence across aggregation algorithms stems from the redundancy in client datasets, where each client’s data mirrors the overall distribution. This redundancy allows the global model to learn similar patterns from any subset of clients, ensuring that convergence remains stable regardless of quantity skew.
8 Spatiotemporal Skew Results
@@
8.1 Synthetic Partitioning Method
The primary constraint in this partition process is that the dataset must contain a categorical variable of space (i.e., places, cities, latitude, longitude, etc.) or time (i.e., hours, months, years, etc.) to use as the partitioning variable. For instance, Figure 5 depicts the distribution of labels along the date of the 5GNTF dataset. In this case, the space variable employed to create the federated data is the flow’s date expressed in year-month-day format (categorical).
We use the St-Dirichlet method from FedArtML [33], which employs the DD to segment the data based on spatial (SP skew) or temporal (TMP skew) categories to distribute the data among federated clients. We assess the level of non-IID data using the HD for spatiotemporal skew (STHD) within the range {0, 0.25, 0.5, 0.75, 0.9}.
8.2 Classification Power
In this subsection, we focus on the simulation results to evaluate various aggregation algorithms and datasets regarding classification accuracy. We pay particular attention to the impact of different levels of spatiotemporal skewness among the clients’ data.
Figure 6: Changes in the models’ accuracy considering different levels of non-IID data measured by STHD for $K=30$.
Highlight 13: The higher the non-IID data level in time and space, the worse the model’s performance.
Table 12: Mean and standard deviation Accuracy for each dataset for CL, FedAvg, Rand, FedProx, Power-Of-Choice, and MOON, considering different levels of non-IID partitioning of the records based on their space (SP) and time (TMP) measured by STHD for $K=30$. Each model has undergone five trials.
@@ 8.2 Classification Power
Number of times that performed the best 0 0 4 3 8
Table 12 depicts the performance for each dataset and aggregation algorithms, using multiple levels of non-IID spatiotemporal skew. It demonstrates that, irrespective of the aggregation algorithm employed, model performance deteriorates when data distribution among clients varies concerning time or space. Figure 6 illustrates the comparison between FedAvg and the centralized model across varying STHD levels, showing the same pattern: model accuracy deteriorates as spatiotemporal non-IID data among clients grows. However, the magnitude of this deterioration differs across datasets. The performance in all datasets drops noticeably once the STHD surpasses 0.75. This phenomenon occurs because increasing the differences among clients’ data based on time and location also raises non-IID data in the clients’ label distributions. This behavior is evident in Table 13, which displays the HD in label distributions at varying levels of STHD among clients. We also concluded in the label skew study section that higher levels of non-IID data among clients’ data distributions negatively impact the performance of the final model.
Table 13: HD among clients’ label distributions at varying levels of non-IID partitioning by time and space
Dataset STHD = 0 STHD = 0.25 STHD = 0.50 STHD = 0.75 STHD = 0.9
Serengeti 0.01 0.09 0.22 0.36 0.53
MHEALTH 0.01 0.01 0.01 0.03 0.07
5GNTF 0.03 0.20 0.29 0.30 0.49
8.3 Convergence
In this subsection, we showcase the results related to the models’ convergence when there are variations in the data concerning time and location, considering the results obtained on the Serengeti, MHEALTH, and 5GNTF datasets.
Table 14: RTA reached for different levels of spatiotemporal non-IID cases as determined by STHD for K = 30.
@@ 8.3 Convergence
5GNTF FedAvg 1 1 1 1 1
Rand 1 1 1 1 1
FedProx 1 1 1 1 1
POC 1 1 1 1 1
MOON 1 1 1 1 1
MHEALTH FedAvg 2 2 2 3 2
Rand 2 2 2 3 2
FedProx 2 2 2 3 2
POC 2 2 2 4 2
MOON 6 7 7 13 14
Highlight 14: Spatial non-IID data shows no consistent impact on convergence rounds; effects vary by dataset dynamics and task difficulty.
According to Table 14, the same level of non-IID data can have a radically different effect depending on the underlying data:
  • Serengeti: As STHD increases from 0 to 0.90, RTA nearly doubles across all aggregation algorithms (e.g., FedAvg: 6 $\rightarrow$ 14; POC: 7 $\rightarrow$ 15). This suggests that data from different sites becomes more heterogeneous, requiring more training rounds for the models to converge.
  • 5GNTF: All aggregation algorithms converge in just one round, despite the STHD. This occurs when the classes are very easy to distinguish and there’s a clear gap between the records of one class and those of the others. In such cases, the task becomes trivial, and the time factor has minimal impact on when convergence is reached.
  • MHEALTH: Across all aggregation algorithms, RTA stays constant—except for MOON, which jumps sharply from 6 to 14 rounds as spatiotemporal non-IID data goes from 0 to 0.90. Since the data come from body-worn sensors and are split by individual subjects, client-specific covariate shifts emerge that particularly undermine representation-based approaches like MOON.
For the MOON algorithm, the gap in RTA between STHD = 0 and STHD = 0.9 varies by dataset—0 rounds for 5GNTF, 8 for MHEALTH, and 11 for Serengeti—while other aggregation algorithms show no such change. This indicates there is not a one-to-one relationship between STHD and convergence speed; instead, factors like task complexity, temporal patterns, and feature diversity shape the outcome, supporting our earlier point that spatial non-IID data impacts convergence inconsistently, depending on dataset dynamics and problem difficulty.
9 General Results
@@
In this section, we provide highlights summarizing the overall results obtained from our experiments, combining the behavior shown before for label, feature, quantity, and spatiotemporal skews.
Highlight 15: Label skew [64, 38] and spatiotemporal skew significantly impact the model’s performance.
Our experimental analysis reveals that not all forms of non-IID data equally degrade FL performance. Label skew and spatiotemporal skew exhibit the most severe impact. Label skew reduces model accuracy by 10–40% compared to the CL baseline. Feature and quantity skews show less significant effects (1-5% accuracy drops). The previous aligns with prior findings [64, 38] that label skew disproportionately harms aggregation, as local models overfit to dominant classes.
Spatiotemporal skew introduces contextual drift (e.g., sensor data varying across locations/times), corrupting the feature space. Similarly to label skew, this cannot be fixed through simple aggregation - our tests show FedAvg suffers 10-12% higher accuracy loss. The global model fails to perform effectively across all contexts because it averages away crucial environmental patterns unique to specific locations or times.
Highlight 16: FedProx performs better under label skew, POC excels in handling quantity skew, and MOON exhibits greater robustness to feature skew, whereas FedAvg and Rand tend to struggle under high non-IID data levels.
@@
Table 15: The number of cases in which each specific algorithm achieved the best performance for each study.
@@
Spatio Temporal Skew 15 0 0 4 3 8
Total best performance 9 3 26 18 32
Table 15 summarizes the cases in which each specific algorithm exhibited the best performance compared to other aggregation algorithms across four skewness types considered in our study. In most cases, the FedProx, POC, and MOON aggregation algorithms achieved the best performance, outperforming the simpler FedAvg and Rand algorithms.
Such a superior performance can be attributed to the specific mechanisms to tackle the non-IID data of each aggregation algorithm. FedProx stabilizes training by regulating the influence of the global model on local clients, POC enhances personalization through loss-based selection, and MOON leverages contrastive learning to improve feature representation. These mechanisms enable better adaptation to diverse client distributions, leading to consistently stronger performance across different skewness types.
Although FedProx, POC, and MOON generally outperform FedAvg and Rand, the performance gains are often marginal. This indicates that while these methods offer improvements in handling non-IID data, they do not fully resolve the challenges posed by non-IID data. The relatively small advantage suggests the need for more effective aggregation algorithms to better adapt to diverse client distributions and enhance model performance in FL scenarios.
Highlight 17: FedProx is more effective on image datasets, while MOON performs better with tabular datasets.
Table 16: The number of cases in which each aggregation algorithm achieved the best performance for each type of dataset
@@
Image 39 8 3 19 9 0
Tabular 49 1 0 7 9 32
Total best performance 9 3 26 18 32
Table 16 examines the best-performing aggregation algorithms from the perspective of the dataset type used for training. It shows that FedProx outperforms all other algorithms on image datasets in ninetheen out of thirty-nine cases. In comparison, MOON generally surpasses other algorithms on tabular datasets in thirty-two out of forty-nine cases.
The effectiveness of FedProx on image datasets and MOON on tabular datasets can be attributed to their distinct optimization strategies. FedProx mitigates client drift by stabilizing updates, which is particularly beneficial for complex, high-dimensional image data. In contrast, MOON’s contrastive learning framework enhances feature representation, making it more suited for tabular data, where feature relationships play a critical role. These differences explain their varying performance across dataset types.
10 Design Insights and Opportunities
We provide some design insights and opportunities, intending to help researchers direct their efforts toward solving the effects of non-IID data.
Quantifying the level of non-IID data.
Several works claim that the non-IID data affects the performance of FL models [44, 45, 31]. Nevertheless, for the first time, we demonstrate that the effect of the non-IID data is not the same under all the levels of heterogeneity (see Figure 3).
Therefore, it is vital to quantify the non-IID data level in FL. This work uses the HD metric to measure the level of non-IID data. However, we encourage researchers to test different metrics, such as JSD [51], EMD [16], and Total Variation distance [7], among others.
More effective methods to tackle high non-IID data.
This work showcases how the state-of-the-art methods to tackle non-IID data (Rand, POC, FedProx, MOON) perform against FedAvg. The conclusion is that no algorithm works better than FedAvg in all the scenarios. Moreover, the better methods do not greatly improve FedAvg under high non-IID scenarios, as their gain is at most two percentage points. FedAvg remains competitive due to its computational efficiency and adequacy for moderately non-IID data, where its simplicity outperforms complex methods like FedProx or MOON that incur tuning overheads. However, in highly non-IID scenarios, adaptive approaches are essential, revealing a core trade-off between simplicity and adaptability. Thus, optimal algorithm selection depends on the specific non-IID data and system constraints in a given FL deployment.
This phenomenon has also been studied and claimed in the scarce work of empirical analysis of non-IID data and methodologies [64, 1, 49, 38]. Therefore, creating methods to alleviate the effect of high levels of non-IID data appropriately is needed to evolve and preserve FL. This aligns with the open problems reported by Kairouz et al. [34].
Focusing on highly unbalanced data.
In our experiments, we claim that the more significant decrease in performance comparing CL and FL occurs in unbalanced datasets since it relates to the challenge of learning from highly skewed and less representative data (see Table 3).
Thus, it is relevant to create solutions to tackle non-IID data by considering the degree of unbalancedness that the labels might have across the clients.